Grade 12Magical Mathematics[Interesting Approach]

Value of (2/3)^infinity or any value less than 1 to the power infinity ......explain …........

milind
12 Years agoGrade 12

1 Answer

bharat bajaj
ApprovedApproved Tutor Answer12 Years ago
Say the number 0.5.
0.5^1 = 0.5
0.5^2 = 0.25
0.5^3 = 0.125
You can see that the value decreases as the power increases of a number less than 1.. Hence, we can say that.. If -1<x<1, then
x^ (inf) tends to 0..
THanks
Bharat Bajaj
askiitians faculty
IIT Delhi